if the price of goods X falls from 55.00 to 54.60 and the demand increases from 1700 units to 2000 units , what would be the price elasticity of demand ?
1
Expert's answer
2019-01-25T11:32:07-0500

If the price of goods X falls from 55.00 to 54.60 and the demand increases from 1700 units to 2000 units, then the price elasticity of demand is:

Ed = (2000 - 1700)/(54.6 - 55)*(54.6 + 55)/(2000 + 1700) = 300/(-0.4)*109.6/3700 = -22.22, so the demand is elastic.